# CrossCheck-Bench: Diagnosing Compositional Failures in Multimodal Conflict Resolution

### v1

@@ Abstract
Figure 2: Overview of CrossCheck-Bench. The benchmark spans three cognitive tiers—L1 Perception, L2 Integration, and L3 Reasoning—grounded in seven atomic capabilities (A1–A7) and eight representative tasks. It offers thousands of product samples, 14,690 adversarial QA pairs, and 22.8 K multimodal clue graphs, curated through 450+ expert hours. The right-hand bar chart contrasts average model accuracy with the human upper bound.
@@ Introduction
Multimodal content in the open world frequently exhibits noisy, unreliable, and even deceptive characteristics. A product page may display a luxury brand logo with a suspiciously low price, or pair an image of a sports shoe with a textual description of formal wear. Humans can instinctively recognize when such visual and textual clues do not align, flagging potential fraud or misrepresentation. This ability to resolve cross-modal conflicts is fundamental for any robust multimodal reasoning system (wang2023aligning; wolf2023fundamental). However, vision-language models (VLMs), which now underpin many content understanding deployments, have not been thoroughly evaluated on their ability to detect and reject contradictions (tan2020detecting). Most existing VLMs (li2023blip; alayrac2022flamingo; liu2023visual) are primarily trained and evaluated on aligned datasets, where vision and language describe the same semantic content. This alignment-centric paradigm encourages cross-modal consistency but overlooks a critical question: can models verify whether multimodal signals are logically compatible? The lack of this capability would pose a tangible risk: models may confidently affirm incompatible clues, producing outputs that are not only inaccurate but also logically inconsistent with the input evidence (li2023evaluating; liu2023trustworthy).
@@
Consequently, there is an urgent need for a benchmark that can effectively evaluate such an ability (ma2023crepe). While existing benchmarks effectively assess tasks like retrieval (yang2022blip; wasserman2025real), description (maaz2023video; yue2024mmmu), and entailment (xie2019visual), few explicitly test whether multimodal inputs are jointly compatible in a structured, diagnostic way. Unlike generation correctness evaluation, validating conflict resolution requires more than local grounding or factual recall—it demands compositional verification across modalities (johnson2017clevr). Specifically, this ability requires models to (1) localize corresponding semantic entities or attributes in each modality and (2) assess whether these aligned clues are logically compatible. However, constructing such data is challenging, as it requires reliable sources for identifying cross-modal contradictions and for synthesising realistic conflict-driven evaluation cases (johnson2017clevr).
@@
To address this gap, we introduce CrossCheck-Bench, the benchmark designed to evaluate and diagnose VLMs’ capacity to resolve multimodal inconsistencies (li2024survey; bai2024hallucination; gunjal2024detecting). As illustrated in Figure 2, we structure the benchmark into a three-level hierarchy reflecting increasing reasoning complexity (yue2024mmmu; lu2024mathvista). L1 (Perceptual Anchoring) evaluates whether the model can extract atomic entities from each modality. L2 (Knowledge Integration) assesses whether the model can compare cross-modal attributes. L3 (Conflict Reasoning) requires the model to detect implicit contradictions that arise from multi-attribute combinations. Each level builds on prior capabilities: from recognizing entities, to comparing attributes, to applying rule-based logic under uncertainty. To enable fine-grained diagnosis, we further decompose the evaluation into seven atomic capabilities spanning Entity Recognition, Attribute Comparison, Numerical Reasoning, Multi-Frame Extraction, etc. By structuring both tasks and skills hierarchically, CrossCheck-Bench reveals how early-stage failures—such as misidentifying logos or misreading text—can propagate upward, leading to confident but flawed high-level inferences. As exemplified by the case study in Figure 1, a single input sample may trigger success at L1 while failing at L2 and L3, illustrating how surface-level perception often masks deeper reasoning collapse. Each QA sample is derived from real-world stimuli (e.g., e-commerce listings, ads, social posts) with injected inconsistencies requiring nontrivial inference.
@@
We evaluated 13 current top performing VLMs, including GPT-4.1 (achiam2023gpt), Gemini-2.5 (comanici2025gemini), Qwen2.5-VL (bai2025qwen2), InternVL3 (zhu2025internvl3), and MiMo-VL (xiaomi2025mimo), on 15k conflict-oriented QA samples spanning all three task levels. Our results reveal a stark performance drop from L1 to L3: while most models succeed at local grounding, nearly all fail at multi-attribute contradiction detection, affirming implausible combinations with high confidence. We further examine whether prompting strategies and lightweight adaptation can enhance model performance on conflict-sensitive tasks. Our findings show that conventional methods such as Chain-of-Thought prompting and visual grounding through annotated regions provide limited benefit and may even reinforce superficial patterns. In contrast, approaches that support iterative reasoning across visual and textual inputs, including lightweight supervised fine-tuning, result in more consistent improvements. Overall, these results highlight a critical bottleneck in the ability of MLLMs to perform robust, integrated reasoning—a key challenge for future research.
@@
Figure 3: Dataset–construction pipeline for CrossCheck-Bench. Stage 1 aggregates multimodal product data (30 + categories, five languages, three modalities) and encodes them as multimodal clue graphs that bind entities to their validated attributes. Stage 2 samples 1–$n$ clues from each graph to programmatically compose hierarchical question–answer pairs that probe the seven atomic capabilities across the three cognitive tiers (L1–L3). Stage 3 applies a three-step quality-control loop: expert review, ensemble-model filtering, and tier-wise difficulty balancing, to ensure answer correctness and uniform task difficulty.
@@
  • We develop a large-scale dataset of 15k QA samples sourced from real-world artifacts, in which real-world contradictions are preserved and programmatically formalized to support systematic evaluation.
@@
Existing benchmarks primarily evaluate VLMs on compositional tasks where modalities reinforce each other. Datasets like VCR (zellers2019recognition), NLVR2 (suhr2019corpus), and SNLI-VE (xie2019visual) assess textual entailment or reasoning with aligned visual support, assuming visual-textual consistency. Recent benchmarks such as MMMU (yue2024mmmu) and MathVista (lu2023mathvista) emphasize complex multimodal reasoning but remain confined to scenarios with concordant inputs. While these efforts demonstrate VLMs’ growing proficiency in integrated understanding, they fail to assess models’ resilience when modalities conflict—the core challenge addressed by our work.
@@
Prior research on multimodal inconsistency has primarily focused on specialized, narrow tasks (elazar2021measuring; tahmasebi2024multimodal). The MMIR benchmark (yan2025multimodal) evaluates inconsistency reasoning in layout-rich artifacts but restricts its scope to predefined error types and lacks granular diagnosis of underlying capability failures. Works like Beyond Appearance (xu2025toward) study modality gaps in specific attributes (e.g., color, shape) but do not scale to compositional real-world conflicts. Crucially, VLM2-Bench (zhang2025vlm2) addresses a distinct challenge: visually linking matching cues across different images (e.g., identifying the same person), which represents an orthogonal problem to resolving contradictory multimodal evidence within unified inputs. CrossCheck-Bench thus fills a critical void by introducing the first hierarchical evaluation framework ($\text{L1}\rightarrow\text{L3}$) to diagnose why models fail at cross-modal conflict resolution.
@@
Efforts to diagnose VLM failures typically dissect atomic capabilities in isolation. SpaCE-10 (gong2025space) decomposes spatial intelligence into 10 atomic skills but does not examine their interplay under conflicting evidence. BEiT-3 (wang2023image) and LLaVA (liu2023visual) analyze modality biases via probing tasks, yet these remain decoupled from real-world inconsistency scenarios. Our benchmark uniquely integrates capability-centric diagnosis with adversarial multimodal conflict: we not only define 7 atomic cross-checking capabilities (e.g., entity grounding, attribute verification) but deliberately stress-test their composition in failure-prone contexts where cues contradict, thereby exposing cascading reasoning breakdowns unseen in prior benchmarks.
@@
To enable fine-grained diagnosis of multimodal inconsistency resolution in Vision-Language Models (VLMs), we introduce CrossCheck-Bench, a factually grounded benchmark comprising 7 atomic capabilities and 15 systematically constructed tasks (Appendix A, B and G). As shown in Figure 2, CrossCheck-Bench follows a hierarchical structure, categorizing tasks into Perception (L1), Integration (L2), and Reasoning (L3) levels, based on the combination of atomic capabilities and task complexity. To ensure the reliability and consistency, the construction of CrossCheck-Bench involves three key stages (see Figure 3): Data Collection via Multimodal Clue Graphs, Hierarchical Tasks Generation, and Quality Verification. In total, over 450 expert-hours were invested to curate 14.69k high-fidelity, semantically grounded QA tasks.
@@
To establish a benchmark that faithfully captures real-world complexity, we compiled a carefully curated dataset of product listings from major e-commerce platforms, selected for their linguistic diversity and market heterogeneity. Each entry includes a minimum of five verified product attributes, a high-resolution product image, and corresponding textual description.
We introduce Multimodal Cue Graphs (MCGs) as a structured representation of factual signals across modalities. Each MCG is a knowledge-graph-like structure composed of quadruples: (entity, modality, attribute, value) (kommineni2024human; yao2025exploring). The construction of MCGs proceeds in three stages. First, entities are extracted via an ensemble of detectors: for images, we combine YOLOv8-L(yi2023small) and GroundingDINO(liu2024grounding) with trademark verification via visual embeddings; for text, we employ Qwen3-8B(yang2025qwen3), fine-tuned for named entity recognition (li2020survey). Second, entity attributes are extracted through rule-based templates and GPT-4o augmentation, incorporating both visual (e.g., OCR, color, shape) and textual (e.g., brand, SKU, price) signals. Third, attribute values are cross-validated between modalities via GPT-4o to identify inconsistencies. Discrepant pairs are flagged and corrected, with a stratified 15% manually audited sample yielding 98.2% accuracy. The final set includes 22.8k MCGs, each containing on average 12.7 verifiable clues, with a semantic consistency rate of 97.3% based on expert auditing.
@@
Question Generation. To generate high-quality and diverse questions aligned with our three-level taxonomy, we adopt a hybrid framework combining template, model, and human strategies. (1) Template-based (L1). For perception-level tasks focusing on atomic skills, we design 45+ rule-based templates operating on MCGs. These templates introduce controlled perturbations to test abilities such as grounding and attribute consistency. We provide examples of the prompt templates in Appendix C. (2) Model-assisted (L2). For integration-level tasks requiring multiple clues, GPT-4o is prompted with structured instructions to locate semantically related MCG facts and produce questions involving cross-modal or temporal reasoning. Human reviewers refine these outputs. (3) Human-authored (L3). For reasoning-level tasks demanding multi-step inference or world knowledge, experts manually craft questions by synthesizing multiple MCG clues. These target compositional challenges such as intention deception or rule violations, ensuring questions are unambiguous and robust against shortcuts. All questions undergo expert validation to remove artifacts and guarantee semantic rigor. The final benchmark contains 14.69k QA pairs balanced across the three cognitive levels. A detailed composition breakdown is shown in Figure 4.
@@
We perform adversarial validation on 40% of samples: three reviewers assess each QA pair for ambiguity or shortcuts, with 12% flagged for revision. Inter-annotator agreement (IAA) on a separate 10% confirms annotation stability. These steps ensure that CrossCheck-Bench is both challenging and reliable for multimodal reasoning evaluation. The detailed verification pipeline is provided in Appendix E.
@@
We evaluate 13 vision-language models (VLMs), including both proprietary and open-source systems capable of processing interleaved image-text inputs. The proprietary models comprise Gemini 2.5 Pro(comanici2025gemini), GPT-4.1 (achiam2023gpt), GPT-4.1mini (achiam2023gpt), and ByteDance-seed-1.6 (seed2025seed1_5vl). The open-source group spans multiple families and scales: Qwen2.5-VL(7B, 32B, 72B)(bai2025qwen2), InternVL3(38B, 78B)(zhu2025internvl3) , Ovis2(8B, 16B, 34B)(lu2024ovis), MiMo-VL-7B(xiaomi2025mimo). All models are evaluated on CrossCheck-Bench using a unified zero-shot QA protocol with standardized prompts. Evaluation employs hybrid scoring: deterministic single-choice questions are assessed via exact match, while open-ended responses are semantically judged by GPT-4o (Appendix D). Open-source models are executed using official implementations on NVIDIA H100 GPUs. Human baseline performance is collected from seven expert annotators following the same QA protocols (Appendix F).
@@
Figure 7: Effect of $M$-CoT. The figure shows the shift in model performance under different prompting interventions. MM-CoT consistently improves both reasoning alignment and overall accuracy across models. Inset: focused comparison among intervention strategies on reasoning-intensive tasks (A5–A7).
@@
We evaluate vision-language models under four strategies: Chain-of-Thought (CoT) prompting(wei2022chain), Set-of-Mark (SoM) visual guidance (yang2023set), a combined CoT+SoM setting, and supervised fine-tuning (CSFT) with 500 curated QA pairs. CoT adds symbolic reasoning scaffolds. SoM uses bounding boxes to guide attention. The combined setup merges both. CSFT tests model adaptability with lightweight supervision.

### v2

@@ Abstract
Figure 2: Overview of CrossCheck-Bench. The benchmark spans three cognitive tiers—L1 Perception, L2 Integration, and L3 Reasoning—grounded in seven atomic capabilities (A1–A7) and eight representative tasks. It offers thousands product samples, 14,690 adversarial QA pairs, and 22.8 K multimodal clue graphs, curated through 450+ expert hours. The right-hand bar chart contrasts average model accuracy with the human upper bound.
@@ Introduction
Multimodal content in the open world frequently exhibits noisy, unreliable, and even deceptive characteristics. A product page may display a luxury brand logo with a suspiciously low price, or pair an image of a sports shoe with a textual description of formal wear. Humans can instinctively recognize when such visual and textual clues do not align, flagging potential fraud or misrepresentation. This ability to resolve cross-modal conflicts is fundamental for any robust multimodal reasoning system (Wang et al. 2023b; Wolf et al. 2023). However, vision-language models (VLMs), which now underpin many content understanding deployments, have not been thoroughly evaluated on their ability to detect and reject contradictions (Tan et al. 2020). Most existing VLMs (Li et al. 2023a; Alayrac et al. 2022; Liu et al. 2023a) are primarily trained and evaluated on aligned datasets, where vision and language describe the same semantic content. This alignment-centric paradigm encourages cross-modal consistency but overlooks a critical question: can models verify whether multimodal signals are logically compatible? The lack of this capability would pose a tangible risk: models may confidently affirm incompatible clues, producing outputs that are not only inaccurate but also logically inconsistent with the input evidence (Li et al. 2023b; Liu et al. 2023b).
@@
Consequently, there is an urgent need for a benchmark that can effectively evaluate such an ability (Ma et al. 2023). While existing benchmarks effectively assess tasks like retrieval (Li et al. 2022; Wasserman et al. 2025), description (Maaz et al. 2023; Yue et al. 2024), and entailment (Xie et al. 2019), few explicitly test whether multimodal inputs are jointly compatible in a structured, diagnostic way. Unlike generation correctness evaluation, validating conflict resolution requires more than local grounding or factual recall—it demands compositional verification across modalities (Johnson et al. 2017). Specifically, this ability requires models to (1) localize corresponding semantic entities or attributes in each modality and (2) assess whether these aligned clues are logically compatible. However, constructing such data is highly challenging, as it depends on reliable data sources to identify contradictory entities across different modalities and create realistic conflict-based evaluation tasks (Johnson et al. 2017).
@@
To address this gap, we introduce CrossCheck-Bench, the benchmark designed to evaluate and diagnose VLMs’ capacity to resolve multimodal inconsistencies (Li et al. 2024; Bai et al. 2024; Gunjal et al. 2024). As illustrated in Figure 2, we structure the benchmark into a three-level hierarchy reflecting increasing reasoning complexity (Yue et al. 2024; Lu et al. 2023). L1 (Perceptual Anchoring) evaluates whether the model can extract atomic entities from each modality. L2 (Knowledge Integration) assesses whether the model can compare cross-modal attributes. L3 (Conflict Reasoning) requires the model to detect implicit contradictions that arise from multi-attribute combinations. Each level builds on prior capabilities: from recognizing entities, to comparing attributes, to applying rule-based logic under uncertainty. To enable fine-grained diagnosis, we further decompose the evaluation into seven atomic capabilities spanning Entity Recognition, Attribute Comparison, Numerical Reasoning, Multi-Frame Extraction, etc. By structuring both tasks and skills hierarchically, CrossCheck-Bench reveals how early-stage failures—such as misidentifying logos or misreading text—can propagate upward, leading to confident but flawed high-level inferences. As exemplified by the case study in Figure 1, a single input sample may trigger success at L1 while failing at L2 and L3, illustrating how surface-level perception often masks deeper reasoning collapse. Each QA sample is derived from real-world stimuli (e.g., e-commerce listings, ads, social posts) with injected inconsistencies requiring nontrivial inference.
@@
We evaluated 13 current top performing VLMs, including GPT-4.1 (Achiam et al. 2023), Gemini-2.5 (Comanici et al. 2025), Qwen2.5-VL (Bai et al. 2025), InternVL3 (Zhu et al. 2025), and MiMo-VL (Xiaomi et al. 2025), on 15k conflict-oriented QA samples spanning all three task levels. Our results reveal a stark performance drop from L1 to L3: while most models succeed at local grounding, nearly all fail at multi-attribute contradiction detection, affirming implausible combinations with high confidence. We further examine whether prompting strategies and lightweight adaptation can enhance model performance on conflict-sensitive tasks. Our findings show that conventional methods such as Chain-of-Thought prompting and visual grounding through annotated regions provide limited benefit and may even reinforce superficial patterns. In contrast, approaches that support iterative reasoning across visual and textual inputs, including lightweight supervised fine-tuning, result in more consistent improvements. Overall, these results highlight a critical bottleneck in the ability of MLLMs to perform robust, integrated reasoning—a key challenge for future research.
@@
Figure 3: Dataset–construction pipeline for CrossCheck-Bench. Stage 1: Clue Encoding. Aggregates multimodal data (30+ categories, 5 languages) into clue graphs binding entities with validated attributes. Stage 2: QA Composition. Samples 1n1-n clues to generate hierarchical QA pairs targeting 7 capabilities across 3 cognitive tiers (L1–L3). Stage 3: Quality Control. Employs a three-step loop (expert review, model filtering, and difficulty balancing) to ensure correctness and task uniformity.
@@
  • We construct a large-scale dataset of 15k QA samples derived from real-world artifacts, with programmatically injected contradictions that require compositional reasoning to resolve.
@@
Existing benchmarks primarily evaluate VLMs on compositional tasks where modalities reinforce each other. Datasets like VCR (Zellers et al. 2019), NLVR2 (Suhr et al. 2019), and SNLI-VE (Xie et al. 2019) assess textual entailment or reasoning with aligned visual support, assuming visual-textual consistency. Recent benchmarks such as MMMU (Yue et al. 2024) and MathVista (Lu et al. 2023) emphasize complex multimodal reasoning but remain confined to scenarios with concordant inputs. While these efforts demonstrate VLMs’ growing proficiency in integrated understanding, they fail to assess models’ resilience when modalities conflict—the core challenge addressed by our work.
@@
Prior research on multimodal inconsistency has primarily focused on specialized, narrow tasks (Elazar et al. 2021; Tahmasebi et al. 2024). The MMIR benchmark (Yan et al. 2025) evaluates inconsistency reasoning in layout-rich artifacts but restricts its scope to predefined error types and lacks granular diagnosis of underlying capability failures. Works like Beyond Appearance (Xu et al. 2025) study modality gaps in specific attributes (e.g., color, shape) but do not scale to compositional real-world conflicts. Crucially, VLM2-Bench (Zhang et al. 2025) addresses a distinct challenge: visually linking matching cues across different images (e.g., identifying the same person), which represents an orthogonal problem to resolving contradictory multimodal evidence within unified inputs. CrossCheck-Bench thus fills a critical void by introducing the first hierarchical evaluation framework ($\text{L1}\rightarrow\text{L3}$) to diagnose why models fail at cross-modal conflict resolution—requiring both structural alignment and logical comparison within unified multimodal inputs.
@@
Efforts to diagnose VLM failures typically dissect atomic capabilities in isolation. SpaCE-10 (Gong et al. 2025) decomposes spatial intelligence into 10 atomic skills but does not examine their interplay under conflicting evidence. BEiT-3 (Wang et al. 2023a) and LLaVA (Liu et al. 2023a) analyze modality biases via probing tasks, yet these remain decoupled from real-world inconsistency scenarios. Our benchmark uniquely integrates capability-centric diagnosis with adversarial multimodal conflict: we not only define 7 atomic cross-checking capabilities (e.g., entity grounding, attribute verification) but deliberately stress-test their composition in failure-prone contexts where cues contradict, thereby exposing cascading reasoning breakdowns unseen in prior benchmarks.
@@
To enable fine-grained diagnosis of multimodal inconsistency resolution in Vision-Language Models (VLMs), we introduce CrossCheck-Bench, a factually grounded benchmark comprising 7 atomic capabilities and 15 systematically constructed tasks. As shown in Figure 2, CrossCheck-Bench follows a hierarchical structure, categorizing tasks into Perception (L1), Integration (L2), and Reasoning (L3) levels, based on the combination of atomic capabilities and task complexity. To ensure the reliability and consistency, the construction of CrossCheck-Bench involves three key stages (see Figure 3): Data Collection via Multimodal Clue Graphs, Hierarchical Tasks Generation, and Quality Verification. In total, over 450 expert-hours were invested to curate 14.69k high-fidelity, semantically grounded QA tasks.
@@
We curated a diverse e-commerce dataset from major platforms, comprising 22.8k listings with $\geq 5$ verified attributes and high-resolution images. To represent factual signals, we define Multimodal Cue Graphs (MCGs) as quadruples: (entity, modality, attribute, value) (Kommineni et al. 2024; Yao et al. 2025). MCG construction proceeds in three stages:
Entity Extraction: We employ an ensemble of YOLOv8-L (Yi et al. 2023), GroundingDINO (Liu et al. 2024), and visual embeddings for image-level detection, alongside fine-tuned Qwen3-8B (Yang et al. 2025) for textual entity recognition. Attribute Extraction: Visual (e.g., OCR, shape) and textual (e.g., brand, price) signals are extracted via rule-based templates and GPT-4o augmentation. Cross-Validation: GPT-4o identifies cross-modal inconsistencies. Discrepant pairs are corrected, with a 15% manual audit yielding 98.2% accuracy.
The final set includes 22.8k MCGs, each containing on average 12.7 verifiable clues, with a semantic consistency rate of 97.3% based on expert auditing.
@@
Question Generation.
To generate diverse questions aligned with our taxonomy, we adopt a hybrid framework combining template, model, and human strategies: (1) Template-based (L1). For atomic tasks, we design 45+ rule-based templates operating on MCGs to test grounding and attribute consistency. (2) Model-assisted (L2). For integration tasks, GPT-4o is prompted with structured instructions to generate questions involving cross-modal or temporal reasoning, refined by human reviewers. (3) Human-authored (L3). For reasoning tasks requiring multi-step inference, experts manually craft questions targeting compositional challenges like intention deception or rule violations. All questions undergo expert validation to ensure semantic rigor. The final benchmark contains 14.69k QA pairs balanced across three levels. A detailed composition breakdown is shown in Figure 4.
@@
We perform adversarial validation on 40% of samples: three reviewers assess each QA pair for ambiguity or shortcuts, with 12% flagged for revision. Inter-annotator agreement (IAA) on a separate 10% confirms annotation stability. These steps ensure that CrossCheck-Bench is both challenging and reliable for multimodal reasoning evaluation.
@@
We evaluate 13 vision-language models (VLMs), including both proprietary and open-source systems capable of processing interleaved image-text inputs. The proprietary models comprise Gemini 2.5 Pro(Comanici et al. 2025), GPT-4.1 (Achiam et al. 2023), GPT-4.1mini (Achiam et al. 2023), and ByteDance-seed-1.6 (Team 2025). The open-source group spans multiple families and scales: Qwen2.5-VL(7B, 32B, 72B)(Bai et al. 2025), InternVL3(38B, 78B)(Zhu et al. 2025) , Ovis2(8B, 16B, 34B)(Lu et al. 2024), MiMo-VL-7B(Xiaomi et al. 2025). All models are evaluated on CrossCheck-Bench using a unified zero-shot QA protocol with standardized prompts. Evaluation employs hybrid scoring: deterministic single-choice questions are assessed via exact match, while open-ended responses are semantically judged by GPT-4o. Open-source models are executed using official implementations on NVIDIA H100 GPUs. Human baseline performance is collected from seven expert annotators following the same QA protocols.
@@
We evaluate vision-language models under four strategies: Chain-of-Thought (CoT) prompting(Wei et al. 2022), Set-of-Mark (SoM) visual guidance (Yang et al. 2023), a combined CoT+SoM setting, and supervised fine-tuning (CSFT) with 500 curated QA pairs. CoT adds symbolic reasoning scaffolds. SoM uses bounding boxes to guide attention. The combined setup merges both. CSFT tests model adaptability with lightweight supervision.
@@
Figure 7: Effect of $M$-CoT. The figure illustrates the performance shift across models with different prompting interventions. MM-CoT consistently enhances reasoning alignment and overall accuracy. Inset: comparison of intervention strategies on reasoning-intensive tasks (A5–A7).